The natural uncertainty in the measurement of speed of an electron in an atom is
estimated to be 2.2x10^4 m/s conceding ideal set-up and error free measurement.
1
Expert's answer
2019-09-12T12:20:25-0400

According to the uncertainty principle find the uncertainty in the measurement of location:


"∆x∆p\geq\frac{\hbar}{2},\\
\space\\
∆x\geq\frac{\hbar}{2m∆v},\\
\space\\
∆x\geq1.65\cdot10^{-8}\text{ m}."